# LLMs for Generating and Evaluating Counterfactuals: A Comprehensive Study

### v1

@@ Abstract
As NLP models become more complex, understanding their decisions becomes more crucial. Counterfactuals (CFs), where minimal changes to inputs flip a model’s prediction, offer a way to explain these models. While Large Language Models (LLMs) have shown remarkable performance in NLP tasks, their efficacy in generating high-quality CFs remains uncertain. This work fills this gap by investigating how well LLMs generate CFs for two NLU tasks. We conduct a comprehensive comparison of several common LLMs, and evaluate their CFs, assessing both intrinsic metrics, and the impact of these CFs on data augmentation. Moreover, we analyze differences between human and LLM-generated CFs, providing insights for future research directions. Our results show that LLMs generate fluent CFs, but struggle to keep the induced changes minimal. Generating CFs for Sentiment Analysis (SA) is less challenging than NLI where LLMs show weaknesses in generating CFs that flip the original label. This also reflects on the data augmentation performance, where we observe a large gap between augmenting with human and LLMs CFs. Furthermore, we evaluate LLMs’ ability to assess CFs in a mislabelled data setting, and show that they have a strong bias towards agreeing with the provided labels. GPT4 is more robust against this bias and its scores correlate well with automatic metrics. Our findings reveal several limitations and point to potential future work directions.
@@ 1 Introduction
In the field of NLP, LLMs have consistently demonstrated remarkable performance across diverse tasks. However, despite significant advancements in counterfactual generation methods, the efficacy of LLMs in producing high-quality CFs remains an open question. Our study bridges this gap by rigorously assessing the inherent capability of LLMs to generate CFs and identifying the most effective ones. We conduct a comprehensive comparison of several common LLMs, spanning different sizes and accessibility levels, evaluating their performance specifically on the counterfactual generation task. Our assessment encompasses standard metrics for counterfactual quality, as well as an in-depth evaluation of language fluency tailored to the context of counterfactual generation. Furthermore, we extend our analysis to data augmentation. We consider generating CFs for two NLU tasks in this study: Sentiment Analysis (SA) and Natural Language Inference (NLI).
Our analysis demonstrates that LLMs are able to generate fluent text. However, they have difficulties in inducing minimal changes. Generating CFs for SA is less challenging than NLI, where LLMs exhibit weaknesses in generating CFs that flip the labels. For data augmentation, CFs from LLMs can be an alternative to human CFs, as they are able to achieve similar performance, while on NLI further improvements are needed. Furthermore, we show a positive correlation between keeping minimal changes and data augmentation performance. This suggests a new direction to generate improved data for augmentation, potentially leading to more efficient augmentation approaches.
We further assess the ability of LLMs to act as evaluators of CFs. By conducting controlled experiments, we show that LLMs have a strong bias to agree with the given labels, even if these are incorrect. Additionally, we show the alignment between GPT4-based evaluation and intrinsic metrics for CFs, indicating that GPT-4 is a reliable evaluator for counterfactual generation. These findings suggest that GPT-4 is a suitable choice of LLM to use when evaluating the quality of generated CFs. A sample of CFs from different LLMs with the corresponding scores is shown in Figure 1. Finally, to facilitate further research, we contribute a new dataset of CFs generated by various LLMs.
@@ 2.1 Intrinsic Evaluation
Given a fixed classifier $f$ and a dataset with $N$ samples $(x_{1},x_{2},\ldots,x_{N})$, $x=(z_{1},z_{2},\ldots,z_{n})$ represents a sequence of $n$ tokens with a ground truth label $y$. A valid counterfactual $x^{\prime}$ should: (1) achieve the desired target label $y^{\prime}$ with (2) minimal changes, and (3) align with likely feature distributions Molnar (2022). Therefore, in this evaluation, we consider the intrinsic properties of Flip Rate, Textual Similarity, and Perplexity that correspond to each criterion, respectively:
@@ 2.3 LLMs for CFs Evaluation
Evaluation with LLMs has been shown to be a valid alternative to human evaluation on various tasks like open-ended story generation and adversarial attacks Chiang and Lee (2023), open-ended questions Zheng et al. (2023), translation Kocmi and Federmann (2023) and natural language generation Liusie et al. (2024). In this work, we examine how well can LLMs evaluate CFs. Detecting mistakes in CFs with LLMs opens the door for iteratively refining CFs Madaan et al. (2023).
@@ 3.2 Generating Countefactuals
  • Generated counterfactuals can be used for data augmentation (an evaluation aspect)
  • Does not require human intervention (e.g., specifying edits or labeling)
  • Does not require additional training in order to make the study computationally feasible
  • The resulting counterfactuals should depend only on the evaluated LLM in order to exclude any other confounding factors
To create the prompt for the LLMs to generate the counterfactual, we combine two techniques: (1) Selecting the closest factual instance to the current instance Liu et al. (2022). Since the provided example has a crucial effect on performance Liu et al. (2022), we select the closest factual/counterfactual pair that has been generated by humans. We use SentenceBERT Reimers and Gurevych (2019) to obtain the latent space representation, and then calculate the distance using cosine similarity from that latent space. (2) Chain-of-Thought (CoT) prompting Wei et al. (2022), showing the necessary steps to generate a counterfactual instance based on a factual one, since it has been shown to help LLMs reason better and provide higher-quality answers. An an overview of the process for generating CFs is depicted in Figure 2.
Specifically, we use the validation set in each dataset as a reference to select the closest example when generating the counterfactual for the train and test sets. After obtaining the pair of closest instances, we apply CoT prompting by defining three steps to generate the counterfactual:
@@
  • Step 3: Replace the words from Step 2 into the original text to obtain the counterfactual instance.
This prompt aligns with the principle of other work Ross et al. (2021); Treviso et al. (2023); Li et al. (2024), which involve identifying significant words that impact the label and altering them to flip the label, thereby generating counterfactual instances. The prompt examples can be found in the Appendix F.
@@ 3.3 LLMs
We compare open-source LLMs with closed-source LLMs. We choose LLAMA-2 Touvron et al. (2023) and Mistral Jiang et al. (2023) as representatives for open-source models, and GPT-3.5 and GPT-4222We use API from https://openai.com/ as a representative for closed-source LLMs. Table 1 summarizes the properties of each LLM.
@@ 4 Results and Discussion
[…]
SA
Expert Humans 51.07 0.16 81.15
Crowd Humans 48.03 0.14 85.66
GPT3.5 49.53 0.16 79.51
GPT4 49.05 0.29 94.03
LLAMA2 7B 46.99 0.64 78.26
LLAMA2 70B 60.46 0.82 96.31
Mistral 7B 48.55 1.06 95.13
Mistral 56B 64.05 1.50 58.55
NLI - premise
Crowd Humans 74.89 0.17 59.13
GPT3.5 71.62 0.15 35.50
GPT4 73.39 0.28 57.12
LLAMA2 7B 70.34 0.36 41.02
LLAMA2 70B 70.74 0.33 58.20
Mistral 7B 78.34 0.36 37.71
Mistral 56B 75.57 0.34 24.32
NLI - hypothesis
Crowd Humans 65.67 0.19 79.75
GPT3.5 51.30 0.19 41.50
GPT4 58.35 0.21 65.88
LLAMA2 7B 59.60 0.28 38.64
LLAMA2 70B 47.69 0.32 48.85
Mistral 7B 39.06 0.46 44.11
Mistral 56B 62.77 0.33 37.79
Table 2: Metrics for intrinsic evaluation. PPL is perpelexity using GPT-2. TS is Levenshtein distance. FR is flip rate with respect to a SOTA classifier.
@@ 4.1 Intrinsic Evaluation
We show the results for the intrinsic evaluation in Table 2. For flip rate, we use SOTA BERT-based models from Morris et al. (2020).
The obtained perplexity values reflect the high fluency of LLMs, some of which are even more fluent than humans. [^3] Note that the shown perplexity values are based on GPT-2 Distance values show that LLMs do not necessarily adhere to conducting minimal changes. One exception here is GPT3.5, whose average distance values resemble that of human-generated CFs. The large distance values for LLMs-generated CFs could be explained by their tendency to overgenerate Guerreiro et al. (2023).
In terms of flip rate, we notice the some LLMs-generated CFs can have a higher flip rate than human-generated CFs on SA, whereas the opposite can can be observed on NLI. Indeed, NLI CFs could be more difficult to generate than CFs for SA, which explains the gap in flip rate between LLMs and humans on the one hand, and GPT-4 and other LLMs on the other hand (this is especially apparent on the NLI - hypothesis test set). This suggests that GPT-4 should be the preferred choice to generate CFs for explaining models behavior. Note that LLAMA2 70B appears to have a higher flip rate on the SA and NLI - premise test sets, but it also has a much lower success rate for generating CFs (cf. Table 11 in Appendix A). We examine the LLMs-generated CFs in more detail in Section 4.4.
We also notice that Mistral 56B, which has not be trained with instruction-tuning or RLHF, produces low quality CFs in terms of the metrics considered here. This highlights the importance of instruction-tuning and RLHF in generating high-quality texts.
This part of the evaluation shows us that LLMs are able to generate fluent CFs, but struggle to induce minimal changes. It also demonstrates that it is challenging to generate NLI CFs that flip the label, whereas generating SA CFs is less difficult.
@@ 4.2 Data Augmentation
We train on both original training data and CFs from different LLMs to see if augmenting the training data leads to an improved performance. For comparison, we conduct data augmentation with human CFs as well. The results for SA and NLI are shown in Table 3 and 4 respectively.
@@
On the crowd CFs and expert CFs test sets for SA including the LLMs-generated CFs lead to improved performance. Perhaps surprisingly, GPT3.5 and LLAMA2 7B provide the most useful CFs for data augmentation. However, augmenting with human CFs works the best. On the original test set, augmenting with CFs does not improve performance. This shows that the gains in performance from data augmentation are visible only if the test set contains challenging examples.
@@
crowd CFs expert CFs original
Original only 92.21 86.27 90.16
Human Crowd 95.90 92.62 89.96
GPT3.5 95.08 91.19 89.96
GPT4 93.65 88.73 90.37
LLAMA2 70B 94.26 90.78 89.55
LLAMA2 7B 95.49 91.19 89.75
Mistral 7B 92.42 88.11 88.93
Mistral 56B 93.85 86.89 84.02
Table 3: Data augmentation results for SA. Classification model is trained on original and LLMs or human-generated CFs. Evaluation is done on 3 test sets: crowd CFs, expert CFs, original with accuracy as a metric.
NLI.
On the crowd premise test set of NLI, which consists of CFs that were generated by changing the premise only, we notice that most of the LLMs-generated CFs help improve the model’s performance by a good margin (> 13 pp). The gap to augmenting with human CFs, however, is still large ($\sim 9$ pp). On the crowd hyothesis test set, most LLMs lead to a lower performance with two exceptions (GPT4 and LLAMA2 70B), where the performance is on par with no data augmentation. However, here too, there is a large gap to human CFs ($\sim 13$ pp). On the original test set, augmenting with LLMs-generated CFs hurts performance, while augmenting with human-generated CFs bring good improvements (5 pp). This shows how high-quality CFs improve the model’s capabilities, and points to a problem with the LLMs-generated CFs for NLI.
@@
crowd Premise crowd Hypothesis original
Original only 39.88 58.50 73.50
Human Crowd 65.38 71.25 78.50
GPT3.5 55.62 51.88 55.00
GPT4 55.88 58.63 70.75
LLAMA2 7B 56.25 51.12 57.50
LLAMA2 70B 53.50 58.13 65.75
Mistral 7B 52.88 48.38 57.75
Mistral 56B 43.75 39.75 45.75
Table 4: Data augmentation results for SA. Classification model is trained on original and LLMs or human-generated CFs with Accuracy metric.
Connection with intrinsic metrics.
We examine the relation between data augmentation performance on the one hand and perplexity and Levenshtein distance on the other hand. The correlation values in Table 5 suggest that CFs with lower distance (to the factual instances) bring more improvements. Indeed, classifiers could be insensitive to small changes Glockner et al. (2018), and having such examples in the training can make classifiers more robust. Having more fluent CFs correlates well with higher performance, but only on SA. SA classifiers could be benefiting from fluent inputs, which are more probable and appear in more contexts. Conversely, the inputs in NLI are short image captions that are not necessarily fluent or grammatical texts.
Compared values SA NLI
Accuracy & --PPL 0.58 -0.05
Accuracy & --TS 0.86 0.51
@@
In summary, only few LLMs produce CFs that come close to human CFs in terms of performance improvements on SA. On NLI, the results are less positive: LLM CFs bring no improvements on two out of three test sets, and the gap to human CFs is still large. CFs with less changes to the factual instances are more beneficial for data augmentation.

### v2

@@ Abstract
As NLP models become more complex, understanding their decisions becomes more crucial. Counterfactuals (CFs), where minimal changes to inputs flip a model’s prediction, offer a way to explain these models. While Large Language Models (LLMs) have shown remarkable performance in NLP tasks, their efficacy in generating high-quality CFs remains uncertain. This work fills this gap by investigating how well LLMs generate CFs for three tasks. We conduct a comprehensive comparison of several common LLMs, and evaluate their CFs, assessing both intrinsic metrics, and the impact of these CFs on data augmentation. Moreover, we analyze differences between human and LLM-generated CFs, providing insights for future research directions. Our results show that LLMs generate fluent CFs, but struggle to keep the induced changes minimal. Generating CFs for Sentiment Analysis (SA) is less challenging than NLI and Hate Speech (HS) where LLMs show weaknesses in generating CFs that flip the original label. This also reflects on the data augmentation performance, where we observe a large gap between augmenting with human and LLM CFs. Furthermore, we evaluate LLMs’ ability to assess CFs in a mislabelled data setting, and show that they have a strong bias towards agreeing with the provided labels. GPT4 is more robust against this bias, but it shows strong preference to its own generations. Our analysis suggests that safety training is causing GPT4 to prefer its generations, since these generations do not contain harmful content. Our findings reveal several limitations and point to potential future work directions.
@@ 1 Introduction
In the field of NLP, LLMs have consistently demonstrated remarkable performance across diverse tasks. However, despite significant advancements in counterfactual generation methods, the efficacy of LLMs in producing high-quality counterfactuals (CFs) remains an open question. Our study bridges this gap by rigorously assessing the inherent capability of LLMs to generate CFs and identifying the most effective ones. We conduct a comprehensive comparison of several common LLMs, spanning different sizes and accessibility levels, evaluating their performance specifically on the counterfactual generation task. Our assessment encompasses standard metrics for CFs quality, as well as an in-depth evaluation of language fluency tailored to the context of counterfactual generation. Furthermore, we extend our analysis to data augmentation. We consider generating CFs for 3 tasks in this study: Sentiment Analysis (SA), Natural Language Inference (NLI), and Hate Speech (HS).
Our analysis demonstrates that LLMs are able to generate fluent text. However, they have difficulties in inducing minimal changes. Generating CFs for SA is less challenging than NLI and HS, where LLMs exhibit weaknesses in generating CFs that flip the labels. For data augmentation, SA CFs from LLMs can be an alternative to human CFs, as they are able to achieve similar performance, while on NLI and HS further improvements are needed. Furthermore, we show a positive correlation between keeping minimal changes and data augmentation performance. This suggests a new direction to generate improved data for augmentation, potentially leading to more efficient augmentation approaches.
We further assess the ability of LLMs to act as evaluators of CFs. We show a sample of CFs from different LLMs with the corresponding scores in Figure 1. By conducting controlled experiments, we show that LLMs have a strong bias to agree with the given labels, even if these labels are incorrect. GPT4 demonstrates strong preference to its own generations. Our analysis suggests that one reason for this preference is safety training, i.e., GPT4 prefers its own generations, because these generations do not contain any harmful content. Finally, to facilitate further research, we contribute a new dataset of CFs generated by various LLMs. [^1] https://github.com/aix-group/llms-for-cfs/
@@ 2.1 Intrinsic Evaluation
Given a fixed classifier $f$ and a dataset with $N$ samples $(x_{1},x_{2},\ldots,x_{N})$, $x=(z_{1},z_{2},\ldots,z_{n})$ represents a sequence of $n$ tokens with a ground truth label $y$. A valid counterfactual $x^{\prime}$ should: (1) achieve the desired target label $y^{\prime}$ with (2) minimal changes, and (3) align with likely feature distributions Molnar (2022). To evaluate these three desiderata, we consider the intrinsic properties of Flip Rate, Textual Similarity, and Perplexity as also suggested in a benchmark for counterfactual evaluation Nguyen et al. (2024):
@@ 2.3 LLMs for CFs Evaluation
Evaluation with LLMs has been shown to be a valid alternative to human evaluation on various tasks like open-ended story generation and adversarial attacks Chiang and Lee (2023), open-ended questions Zheng et al. (2023), translation Kocmi and Federmann (2023) and natural language generation Liusie et al. (2024). In this work, we examine how well LLMs can evaluate CFs. Detecting mistakes in CFs with LLMs opens the door for iteratively refining CFs Madaan et al. (2023).
@@ 3.1 Data
Hate Speech (HS).
We use the dataset from Vidgen et al. (2021), which includes human CFs.
@@ 3.2 Generating Countefactuals
  • Generated CFs can be used for data augmentation (an evaluation aspect)
  • Generating CFs does not require human intervention (e.g., specifying edits or labeling)
  • Generating CFs does not require additional training in order to make the study computationally feasible
  • The resulting CFs should depend only on the evaluated LLM in order to exclude any other confounding factors
To create the prompt for the LLMs to generate CFs, we combine two techniques: (1) Selecting the closest factual instance to the current instance Liu et al. (2022). Since the provided example has a crucial effect on performance Liu et al. (2022), we select the closest factual/counterfactual pair that has been generated by humans. We use SentenceBERT Reimers and Gurevych (2019) to obtain the latent space representation, and then calculate the distance using cosine similarity from that latent space. (2) Chain-of-Thought (CoT) prompting Wei et al. (2022), showing the necessary steps to generate a counterfactual instance based on a factual one, since it has been shown to help LLMs reason better and provide higher-quality answers. An overview of the process for generating CFs is depicted in Figure 2.
Specifically, we use the validation set in each dataset as a reference to select the closest example when generating CFs for the train and test sets. After obtaining the pair of closest instances, we apply CoT prompting by defining three steps to generate the counterfactual:
@@
  • Step 3: Replace the words from Step 2 in the original text to obtain the counterfactual instance.
This prompt aligns with other work Ross et al. (2021); Treviso et al. (2023); Li et al. (2024), which involve identifying significant words that impact the label and altering them to flip the label, thereby generating counterfactual instances. The prompt examples can be found in the Appendix E.
@@ 3.3 LLMs
We compare open-source LLMs with closed-source LLMs. We choose LLAMA-2 Touvron et al. (2023) and Mistral Jiang et al. (2023) as representatives for open-source models, and GPT-3.5 and GPT-4333We use API from https://openai.com/ as representatives for closed-source LLMs. Table 1 summarizes the properties of each LLM.
@@ 4.1 Intrinsic Evaluation
We show the results for the intrinsic evaluation in Table 6. For flip rate, we use SOTA BERT-based models from Morris et al. (2020) (SA and NLI) and Vidgen et al. (2021) (HS).
The obtained perplexity values reflect the high fluency of LLMs, some of which are even more fluent than humans. [^4] Note that the shown perplexity values are based on GPT-2The perplexity of HS is significantly higher than that of other datasets due to the informal nature of tweets, where users often use slang, uncommon words, or elongated words for emphasis. Distance values show that LLMs do not necessarily adhere to conducting minimal changes. One exception here is GPT3.5, whose average distance values resemble that of human-generated CFs. The large distance values for LLM-generated CFs could be explained by their tendency to overgenerate Guerreiro et al. (2023).
In terms of flip rate, we notice that some LLM-generated CFs can have a higher flip rate than human-generated CFs on SA, whereas the opposite can be observed on NLI. Meanwhile, LLM-generated CFs can reach moderate FR in HS. NLI CFs could be more difficult to generate than SA and HS CFs, which explains the gap in flip rate between LLMs and humans on the one hand, and GPT4 and other LLMs on the other hand (this is especially apparent on the NLI - hypothesis). This suggests that GPT4 should be the preferred choice to generate CFs for explaining a model’s behavior. Furthermore, across all datasets, LLMs struggle to flip the label while keeping the changes minimal, i.e., they often need to make many modifications to flip the label. We examine the LLM-generated CFs in more detail in Section 4.4.
This part of the evaluation shows us that LLMs are able to generate fluent CFs, but struggle to induce minimal changes. It also demonstrates that it is challenging to generate NLI and HS CFs that flip the label, whereas generating SA CFs is less difficult.
SA NLI - premise NLI - hypothesis Hate Speech
PPL $\downarrow$ TS $\downarrow$ FR $\uparrow$ PPL $\downarrow$ TS $\downarrow$ FR $\uparrow$ PPL $\downarrow$ TS $\downarrow$ FR $\uparrow$ PPL $\downarrow$ TS $\downarrow$ FR $\uparrow$
Human Experts 51.07 0.16 81.15 - - - - - - - - -
Human Crowd 48.03 0.14 85.66 74.89 0.17 59.13 65.67 0.19 79.75 229.05 0.31 87.39
GPT3.5 49.53 0.16 79.51 71.62 0.15 35.50 51.30 0.19 41.50 235.52 0.16 54.05
GPT4 49.05 0.29 94.03 73.39 0.28 57.12 58.35 0.21 65.88 209.49 0.49 76.54
LLAMA2 7B 46.99 0.64 78.26 70.34 0.36 41.02 59.60 0.28 38.64 - - -
LLAMA2 70B 33.88 1.37 93.48 63.17 0.21 41.07 58.54 0.23 46.62 - - -
Mistral 7B 48.55 1.06 95.13 78.34 0.36 37.71 39.06 0.46 44.11 365.15 0.69 67.41
Mistral 56B 35.63 0.57 95.45 65.37 0.23 27.46 57.65 0.21 31.55 401.63 0.56 70.30
Table 2: Metrics for intrinsic evaluation. PPL is perpelexity using GPT-2. TS is Levenshtein distance. FR is flip rate with respect to a SOTA classifier. [^6] LLAMA-2 is unable to generate counterfactuals for HS due to its safety mechanism.
@@ 4.2 Data Augmentation
We train on both original training data and CFs from different LLMs to see if augmenting the training data leads to an improved performance. For comparison, we conduct data augmentation with human CFs as well. The results for SA, NLI and HS are shown in Table 3,  4 and 5 respectively.
@@
On the crowd CFs and expert CFs test sets for SA including LLM-generated CFs lead to improved performance. LLAMA2 7B provide the most useful CFs for data augmentation, but other LLMs perform similarly. However, augmenting with human CFs works the best. On the original test set, augmenting with CFs does not improve performance. This shows that the gains in performance from data augmentation are visible only if the test set contains challenging examples.
NLI.
On the crowd premise test set of NLI, which consists of CFs that were generated by changing the premise only, we notice that most of the LLM-generated CFs help improve the model’s performance by a good margin (> 7 pp). The gap to augmenting with human CFs, however, is still large ($\sim 9$ pp). On the crowd hyothesis test set, all LLMs lead to a lower performance. Here too, there is a large gap to human CFs ($\sim 16$ pp). On the original test set, augmenting with LLM-generated CFs hurts performance, while augmenting with human-generated CFs bring good improvements ($\sim 5$ pp). This shows how high-quality human CFs improve the model’s capabilities, and points to a problem with LLM-generated CFs for NLI.
HS.
Training with LLM-generated CFs does not bring substantial improvements on the CFs and the original test sets. Conversely, training with human CFs leads to significant improvements on both test sets. On this task too, LLM-generated CFs fall short of human CFs, indicating that there remains significant room for improvement.
Connection with intrinsic metrics.
We examine the relation between data augmentation performance on the one hand and perplexity and Levenshtein distance on the other hand. The correlation values in Table 6 suggest that CFs with lower distance (to the factual instances) bring more improvements. Indeed, classifiers could be insensitive to small changes Glockner et al. (2018), and having such examples in the training can make classifiers more robust. The negative correlation between accuracy and perplexity suggests that more fluent CFs are less effective in improving the classifier’s performance. This indicates that classifiers primarily focus on the content rather than grammatical structure or coherence, especially in NLI tasks where the (factual) instances are mere image captions that are not necessarily fluent or grammatical texts
In summary, most LLMs produce CFs that come close to human CFs in terms of data augmentation performance on SA. On NLI and HS, the results are less positive: LLM-generated CFs bring no improvements in most cases, and the gap to human CFs is still large. CFs with less changes to the factual instances are more beneficial for data augmentation.
@@
Crowd CFs Expert CFs Orig.
Original only 91.68 ± 1.07 86.31 ± 1.62 90.20 ± 0.67
Human Crowd 95.94 ± 0.37 92.01 ± 1.09 89.63 ± 0.85
GPT3.5 94.55 ± 0.96 89.88 ± 1.47 89.30 ± 0.51
GPT4 93.52 ± 0.89 89.10 ± 0.76 89.88 ± 0.57
LLAMA2 7B 95.29 ± 0.72 90.37 ± 1.57 88.89 ± 1.35
LLAMA2 70B 94.18 ± 0.27 88.89 ± 1.02 89.39 ± 0.44
Mistral 7B 93.93 ± 0.62 88.61 ± 1.68 89.22 ± 0.72
Mistral 56B 93.40 ± 1.02 88.20 ± 0.79 89.84 ± 0.79
Table 3: Data augmentation results for SA. Classification model is trained on original and LLMs or human-generated CFs with Accuracy as a metric.
@@
crowd Premise crowd Hypothesis Orig.
Original only 43.60 ± 3.87 59.75 ± 3.06 71.85 ± 1.33
Human Crowd 63.42 ± 2.74 70.53 ± 1.02 76.65 ± 2.04
GPT3.5 54.42 ± 1.86 49.68 ± 2.64 53.00 ± 2.61
GPT4 53.10 ± 1.85 54.50 ± 1.28 63.50 ± 1.31
LLAMA2 7B 52.85 ± 1.29 49.45 ± 2.03 58.15 ± 2.23
LLAMA2 70B 54.58 ± 3.69 49.02 ± 2.96 58.05 ± 0.78
Mistral 7B 51.05 ± 2.89 46.52 ± 2.51 58.50 ± 2.50
Mistral 56B 51.35 ± 1.79 45.45 ± 1.07 48.65 ± 1.88
Table 4: Data augmentation results for NLI. Classification model is trained on original and LLMs or human-generated CFs with Accuracy metric.
Test Data
CFs Orig.
Original only 83.27 ± 2.66 70.28 ± 0.60
Human 94.27 ± 0.20 94.30 ± 0.14
GPT3.5 81.00 ± 2.87 70.29 ± 0.96
GPT4 86.00 ± 3.20 69.33 ± 0.49
Mistral 7B 84.32 ± 2.52 69.90 ± 0.75
Mistral 56B 82.86 ± 1.78 68.58 ± 0.97
Table 5: Data augmentation results for Hate Speech. Classification model is trained on original and LLMs or human-generated CFs with accuracy as a metric.
Compared values SA NLI HS
Accuracy & --PPL -0.26 -0.56 -0.10
Accuracy & --TS 0.49 0.52 0.60